# InterFormer: Towards Effective Heterogeneous Interaction Learning for Click-Through Rate Prediction

### v2

@@ 6. Conclusion
In this paper, we propose a novel building block for heterogeneous interaction learning in CTR prediction named InterFormer. The key idea is an interleaving learning style between different data modes to achieve effective inter-mode interaction, as well as selective information aggregation. In general, A global arch and a sequence arch are adopted to achieve behavior-aware interaction learning and context-aware sequence modeling, respectively. A bridging arch is further proposed to for effective information selection and summarization. InterFormer achieves consistent outperformance on benchmark datasets, and 0.15% improvement on the NE of CTR prediction on internal large-scale dataset.
@@ A.1. Algorithm
0:  Global feat $\mathbf{X}^{(0)}$, sequence feat $\mathbf{S}^{(0)}$, number of layers $L$;
@@
2:  Compute global summarization $\mathbf{X}_{\text{sum}}^{(1)}$ via Eq. (10);
@@
5:     Bridging Arch: compute global summarization $\mathbf{X}^{(l)}_{\text{sum}}$ and sequence summarization $\mathbf{S}^{(l)}_{\text{sum}}$ by Eqs. (10) and (11);
6:     Global Arch: compute global embeddings $\mathbf{X}^{(l+1)}$ given $\mathbf{X}^{(l)}$ and $\mathbf{S}^{(l)}_{\text{sum}}$ by Eq. (7);
@@ A.2.1. Linear Compressed Embedding
Given numerous global features, it is essential to compress embeddings to a more manageable size. In general, given $N$ $d$-dimensional features $\mathbf{X}\in\mathbb{R}^{d\times N}$, LCE is a linear transformation on the sample dimension $\mathbf{W}\in\mathbb{R}^{N\times M}$, such that the linear transformation $\mathbf{XW}\in\mathbb{R}^{d\times M}$ serves as the compressed embedding with $M$ features. Together with self-gating modules, LCE selectively compress and denoise numerous embeddings, benefiting the model in both efficiency and effectiveness.
@@ A.2.2. Personalized FFN
PFFN learns interactions between global and sequence embeddings, whose key idea is to apply an FFN on the sequence embeddings with weight learned based on the summarized global embeddings. Given global summarization $\mathbf{X}_{\text{sum}}\in\mathbb{R}^{d\times n_{\text{sum}}}$ and sequence embedding $\mathbf{S}\in\mathbb{R}^{d\times T}$, PFFN first learns the transformation weight $\mathbf{W}_{\text{PFFN}}=\mathbf{X}_{\text{sum}}\mathbf{W}\in\mathbb{R}^{d\times d}$, which is further applied on the feature dimension of sequence embeddings, i.e., $\mathbf{W}_{\text{PFFN}}\mathbf{S}$. In contrast to the vanilla FFN in Transformer that may not add too much value to subsequent sequence modeling, PFFN integrate side information from other data modes. In fact, PFFN is essentially computing enhanced dot-products based interaction between sequence embeddings and summarized global embeddings that brings much more value for the following sequence modeling.
@@ B.1. Datasets
  • AmazonElectronics (He and McAuley, 2016b) contains product reviews and metadata from Amazon with 192,403 users, 63,001 goods, 801 categories and 1,689,188 samples. Global features include user id, item id and item category. Sequence features include interacted items and corresponding categories with length of 100. We follow the public split with 2.60M samples for training and 0.38M samples for testing.
  • TaobaoAds (Tianchi, [n. d.]) contains 8 days of ad click-through data on Taobao (26 million records) randomly sampled from 1,140,000 users. Global features include item-related features, e.g., ad id, category, and price, and user-related features, e.g., user id, gender, and age. Sequence features include interacted items’ brands and categories, and user behaviors with length of 50. We follow the public split with 22.0M samples for training and 3.1M samples for testing.
  • KuaiVideo (Li et al., 2019) contains 10,000 users and their 3,239,534 interacted micro-videos. Global features include user id, video id, and visual embeddings of videos. Sequence features include different behaviors, e.g., click, like, and not click, with length of 100. We follow the public split with 10.9M samples for training and 2.7M samples for testing.
  • Internal contains 70B entries in total, with hundreds of global features describing users and items, and 10 sequence features of length 200 to 1,000.

### v3

@@ 6. Conclusion
In this paper, we propose a novel building block for heterogeneous interaction learning in CTR prediction named InterFormer. The key idea is an interleaving learning style between different data modes to achieve effective inter-mode interaction, as well as selective information aggregation. In general, An Interaction Arch and a Sequence Arch are adopted to achieve behavior-aware interaction learning and context-aware sequence modeling, respectively. A Cross Arch is further proposed to for effective information selection and summarization. InterFormer achieves consistent outperformance on benchmark datasets, and 0.15% improvement on the NE of CTR prediction on internal large-scale dataset. Pilot launches in key Ads models have yielded a 0.6% improvement in topline metrics and we expect great impact to be delivered.
Appendix
@@ A.1. Algorithm
0:  Non-sequence feature $\mathbf{X}^{(0)}$, sequence feature $\mathbf{S}^{(0)}$, number of layers $L$;
@@
2:  Compute non-sequence summarization $\mathbf{X}_{\text{sum}}^{(1)}$ via Eq. (10);
@@
5:     Cross Arch: compute non-sequence summarization $\mathbf{X}^{(l)}_{\text{sum}}$ and sequence summarization $\mathbf{S}^{(l)}_{\text{sum}}$ by Eqs. (10) and (11);
6:     Interaction Arch: compute non-sequence embeddings $\mathbf{X}^{(l+1)}$ given $\mathbf{X}^{(l)}$ and $\mathbf{S}^{(l)}_{\text{sum}}$ by Eq. (7);
@@ A.2.1. Linear Compressed Embedding
Given numerous non-sequence features, it is essential to compress embeddings to a manageable size. In general, given $N$ $d$-dimensional features $\mathbf{X}\in\mathbb{R}^{d\times N}$, LCE is a linear transformation on the sample dimension $\mathbf{W}\in\mathbb{R}^{N\times M}$, such that the linear transformation $\mathbf{XW}\in\mathbb{R}^{d\times M}$ serves as the compressed embedding with $M$ features. Together with self-gating modules, LCE selectively compress and denoise numerous embeddings, benefiting the model in both efficiency and effectiveness.
@@ A.2.2. Personalized FFN
PFFN learns interactions between non-sequence and sequence embeddings, whose key idea is to apply an FFN on the sequence embeddings with weight learned based on the summarized non-sequence embeddings. Given non-sequence summarization $\mathbf{X}_{\text{sum}}\in\mathbb{R}^{d\times n_{\text{sum}}}$ and sequence embedding $\mathbf{S}\in\mathbb{R}^{d\times T}$, PFFN first learns the transformation weight $\mathbf{W}_{\text{PFFN}}=\mathbf{X}_{\text{sum}}\mathbf{W}\in\mathbb{R}^{d\times d}$, which is further applied on the feature dimension of sequence embeddings, i.e., $\mathbf{W}_{\text{PFFN}}\mathbf{S}$. In contrast to the vanilla FFN in Transformer that may not add too much value to subsequent sequence modeling, PFFN integrate side information from other data modes. In fact, PFFN is essentially computing enhanced dot-products based interaction between sequence embeddings and summarized non-sequence embeddings that brings much more value for the following sequence modeling.
@@ B.1. Datasets
  • AmazonElectronics (He and McAuley, 2016b) contains product reviews and metadata from Amazon with 192,403 users, 63,001 goods, 801 categories and 1,689,188 samples. Non-sequence features include user id, item id and item category. Sequence features include interacted items and corresponding categories with length of 100. We follow the public split with 2.60M samples for training and 0.38M samples for testing.
  • TaobaoAds (Tianchi, [n. d.]) contains 8 days of ad click-through data on Taobao (26 million records) randomly sampled from 1,140,000 users. Non-sequence features include item-related features, e.g., ad id, category, and price, and user-related features, e.g., user id, gender, and age. Sequence features include interacted items’ brands and categories, and user behaviors with length of 50. We follow the public split with 22.0M samples for training and 3.1M samples for testing.
  • KuaiVideo (Li et al., 2019) contains 10,000 users and their 3,239,534 interacted micro-videos. Non-sequence features include user id, video id, and visual embeddings of videos. Sequence features include different behaviors, e.g., click, like, and not click, with length of 100. We follow the public split with 10.9M samples for training and 2.7M samples for testing.
  • Internal contains 70B entries in total, with hundreds of non-sequence features describing users and items, and 10 sequence features of length 200 to 1,000.